OpenStudy (anonymous):

Write an equation in slope intercept form given the graph below. x -2 -1 0 1 y 2 -3 -8 -13 Answer y = -5x – 8 y = -1/5x + 8 y = -3x + 8 y = 1/3x – 8

OpenStudy (anonymous):

D or C?

OpenStudy (anonymous):

its either a or b since the y intercept is -8 (given the point (0,8)) now just find the slope between two points, easiest by plugging in the y intercept and a set of points into y=mx+b:\[2=m(-2)-8\]\[2=-2m-8\]\[10=-2m\]\[m=-5\]so in conclusion the equation is:\[y=-5x-8\]option A :)